Circular No. 5752

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 K. W. Weiler, Naval Research Laboratory; R. A. Sramek, National Ra-

dio Astronomy Observatory; S. D. Van Dyk, Naval Research Laboratory; and N. Panagia, Space Telescope Science Institute, report: “New observations of SN 1993J made on Apr. 2.30 UT with the Very Large Array in ‘B’ configuration show a possible detection at wavelength 1.3 cm (22.5 GHz) of a source with a flux density of 0.8 ± 0.2 mJy at α = 9^{h}51^{m}19^{s}.032, δ = +69^{o}15′26′′.47 (equinox 1950.0), near the reported optical position of the supernova (IAUC 5731, 5737). It should be emphasized that this possible detection is barely significant and must be confirmed. Observations at 2.0 cm (14.9 GHz) show no detection to a limit (3σ) of < 0.33 mJy. New observations at 3.6 cm (8.4 GHz), combined with previous observations at the same wavelength as reported on IAUC 5743, still show no detection to a very low limit (3σ) of < 0.080 mJy. Radio monitoring is continuing at these and other wavelengths. New measurements of the compact radio source in the nucleus of NGC 3031 provide the following improved position: α = 9^{h}51^{m}27^{s}.321, δ = +69^{o}18′08′′.20 (equinox 1950.0; ± 0′′.2 in each coordinate). The measured peak flux densities of the nucleus are 59 mJy at 3.6 cm and 12 mJy at 20 cm.”

COMET SHOEMAKER-LEVY (1993e) B. A. McLeod and K. K. McLeod, Steward Observatory, report obser-

vations made at 2.2 μm with the Steward 1.6-m telescope on Mt. Bigelow with a NICMOS3 detector. The 22-min mosaic centered on Apr. 2.37 UT showed the comet to be 60′′ long in p.a. 75^{o}-255^{o}. The integration showed emission extending 10′′ to the northwest at the eastern end of the bar and 20′′ to the northwest at the western end. The integrated K magnitude within a 90′′ × 55′′ region was 12.0 ± 0.15.

SUPERNOVA 1993G in NGC 3690 J. Stull, Alfred University Observatory, provides the following informa-

tion, obtained using a 0.8-m reflector and unfiltered ST-6 CCD. Assuming a magnitude of 16.6 on Mar. 12.09 UT to calibrate the western core of the galaxy as the comparison object, it followed that the magnitudes of the supernova were 16.7 and 17.0 on Mar. 19.16 and 26.16, respectively.

1993 April 6 Daniel W. E. Green